OpenStudy (anonymous):

SOMEONE PLEASE HELP Find t* for a 95% confidence intervals if the sample size is 17. Question 1 options: t* = 2.09 t* = 1.75 t* = 2.12 t* = 1.70

jimthompson5910 (jim_thompson5910):

use a table like this https://people.richland.edu/james/lecture/m170/tbl-t.html which is normally found in the appendix of your stats textbook

jimthompson5910 (jim_thompson5910):

look at the `95%` confidence level column and locate the df = n-1 = 17-1 = 16 row

OpenStudy (anonymous):

@jim_thompson5910 I got 2.12, is that right?

OpenStudy (nikato):

^yes, that is correct

jimthompson5910 (jim_thompson5910):

yep that's correct